Arizona LBTI Commissioning - 16 October 2010 UT

This wiki page includes John's notes on the center bent Gregorian focal station commissioning aspects of LBTI Commissioning. This represents only a small fraction of the total activity.

Observers: JHill, PHinz, Vidhya, Bill
Telescope Operators: SAllanson, GBechetti
Instrument Support: JMorris
Telescope Support: JUrban

Summary

Worked on Adaptive Optics interfaces in the early evening. We found the routine fringes again around midnight. Closed at 2AM for clouds.

Preparations

See the low priority issue track about laptops not being able to access the weather URLs.

John added side flags to ptspiral.py so it is binocular PCS compatible. (ptspiral_left and ptspiral_right should be obsolete, although I haven't deleted them yet.)

TeamAO helped Vidhya,Oli,Paul get the SX Acquisition Camera working.

Michele patched AO5 PCS to get rid of the phantom rotator problem from last night.

Thunderstorms over Tucson in the late afternoon, but still dry here on the mountain (so far).

John updated the Zglobal focus terms for both primary mirrors to match the values we observed last night. (No coma corrections or tip-tilts have been included yet.)

DaveT gave us an improved version of LBTtools with more side awareness.

Details

Adaptive Optics Tests

01:00 Simone, Lorenzo, Marco and Fernando are helping Vidhya and friends get the adaptive correction running for LBTI. Since the retroreflector was on during the day, they have kept it on into the early evening for these tests.

03:00 Sky is clearing off. Still some heavy stuff off to the southeast.

04:00 Removing the retroreflector from DX M2, and debugging why the slopes are not arriving from the LBTI WFS BCU to the M2 BCU.

Open

04:28 Open to sky with a few clouds. T=5.6 degC, D=-3.3 degC Wind light.

04:34 To zenith to study the BCU fiber configuration.

05:13 Humidity is up to 90%. Mirrors and telescope are slowly cooling.

05:30 Steve sends a test preset in track mode. It fails after spending 5 minutes waiting for Azimuth & Rotator (azimuth was on-target).

Preemptively set IE=-50 CA=+37.

Pointing at a real target star

05:40 Vidhya sends a track preset that succeeded AND put a star on MIRAC. Phil is impressed with our dead aim. The elevation is 76. Hers was BOTH & IDLE, while Steve's was LEFT & POSITION.

05:43 Phil's offset command works, although perhaps not in the direction he was expecting.

05:49 Send DX Tip=-16 Tilt=-43 as M1M2 Later, this was found to be the wrong direction.

05:50 Send SX tip=5 as M1M2 and then back, but the star didn't return to exactly the same position. Probably because of a collimation update.

05:54 Work pauses on sky for clouds drifting by. Humidity is down to 84%. Switch to 2 micron imaging.

06:05 -1000 Z7 to SX Primary (image was >)

06:07 SX M1 Zglobal=0.22

SX is on MIRAC at IE=-54 CA=+39

06:18 Put DX back to zero RX-RY, Put SX to 60 RX to move it out of the way.

Ran the full spiral of 10 semi-loops at 7 arcsec pitch without finding a star on the DX side. In the meantime, Phil and Vidhya are puzzling over the peculiar pattern on the DX acq camera.

06:26 Move to zenith to check the DX dichroic. It was OK.

06:37 Resend preset. for SX IE=-52 CA=+41 Zglobal=+0.16

Switch to pupil mode to check that view.

06:45 Rotate DX M3 to verify which pupil is which. The upper pupil on the MIRAC display is DX as expected.

Now we think we have two problems. The DX pupil looks fine, but has no star image. The SX pupil looks really bad (like something is covering the aperture.) Oops, we thought wrong. The extra signal on the SX side was the actual starlight! So we are back to the single mystery of why we don't see the star in DX.

06:56 Back to imaging mode. DX M3 is back in position after restarting.

On the first spiral above, the M1M2 offset on DX put the star 80 arcsec from the expected location. This second time John had removed the M1M2 offset that was sign reversed, but he also flipped the focus correction from the correct one entered in the afternoon. That's why we had so much trouble.

07:13 Found quite aberrated DX image at IE=-56 CA=2

DX Z6 -7000
+1300 Z7 -1500 Z8

Find DX with Tilt=-39 on M2, and then switched that to +39 on M1M2 Ry with an apparent sign reversal (keep that position)

Had to Put -1200 Z8 on DX M2 as DX M1 ran out of X-range

07:35 SX Zglobal 0.00, DX Zglobal -0.9

Fringe work at 10 microns

They sky has mostly clear but the occasional bit of fluff drifts by. Seeing is getting down toward 1 arcsec.

07:58 The fringes are found again. Steve observes that the room is substantially more subdued that last night.

08:05 Phil was sending offsets to move the star out of the field to measure the background. They were being sent as RIGHT figuring that it doesn't make any difference when you are in monocular track mode. Indeed both stars moved away correctly when the mount moved. However, the offset command was only recollimating the RIGHT mirror (as advertised). This type of offset needs to be sent as BOTH.

08:30 Clouds have rolled over us. We can still see Orion, but that is about all.

Preset to Alpha Tau, but we don't see anything because of clouds.

Close

09:06 Close for thickening clouds.

Operator Training Exercise

We found some things wrong that weren't booby traps laid by John.

09:15 Found SX M3 out of its collimation position. Why? We recollimated and it was OK. Steve says it was OK at the beginning of the night. Did some offset go to the wrong OPE?

09:40 Found DX M3 Swing Arm not showing its deployed limit switch, even though it is deployed. We retracted and then redeployed the swing arm and the switch worked OK.

The DRAIN PIPE HEAT TRACE was found on when it should be off for observing (and summer). We wonder how long it has been on?

-- JohnHill - 16 Oct 2010
Topic attachments
I Attachment Action Size Date Who Comment
pmcc.dx.thermal.20101016.jpgjpg pmcc.dx.thermal.20101016.jpg manage 555 K 18 Oct 2010 - 10:58 JohnHill Plot of DX Primary Mirror Temperatures
pmcc.sx.thermal.20101016.jpgjpg pmcc.sx.thermal.20101016.jpg manage 570 K 18 Oct 2010 - 10:59 JohnHill Plot of SX Primary Mirror Temperatures
Topic revision: r6 - 18 Oct 2010, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback